# EVM-QuestBench: An Execution-Grounded Benchmark for Natural-Language Transaction Code Generation

### v1

@@ Abstract
Large language models are increasingly applied to various development scenarios. However, in on-chain transaction scenarios, even a minor error can cause irreversible loss for users. Existing evaluations often overlook execution accuracy and safety. We introduce EVM-QuestBench, an execution-grounded benchmark for natural-language transaction-script generation on EVM-compatible chains. The benchmark employs dynamic evaluation: instructions are sampled from template pools, numeric parameters are drawn from predefined intervals, and validators verify outcomes against these instantiated values. EVM-QuestBench contains 107 tasks (62 atomic, 45 composite). Its modular architecture enables rapid task development. The runner executes scripts on a forked EVM chain with snapshot isolation; composite tasks apply step-efficiency decay. We evaluate 20 models and find large performance gaps, with split scores revealing persistent asymmetry between single-action precision and multi-step workflow completion. Code: https://anonymous.4open.science/r/bsc_quest_bench-A9CF/.
EVM-QuestBench: An Execution-Grounded Benchmark for Natural-Language Transaction Code Generation
Pei Yang1thanks: Equal contribution.   wanyi Chen1,211footnotemark: 1   Ke Wang1   Lynn Ai1   Eric Yang1   TIANYU SHI1 1Gradient 2Soochow University
@@ 1 Introduction
Large language models (LLMs) are increasingly being used to control software and tools (OpenAI, 2023; Anil and others, 2023). Leveraging LLMs for code generation and blockchain transactions is becoming commonplace, but this introduces significant financial risks. Even a small error, such as an incorrect address, unit, or deadline, can result in irreversible losses.
Benchmarks for code understanding and generation need to cover a broad range of programming tasks and reasoning capabilities (Chen et al., 2021; Lu et al., 2021). Many evaluations still rely on lexical overlap metrics such as BLEU or CodeBLEU (Papineni et al., 2002; Ren and others, 2020). These metrics can reward outputs that appear similar to references but fail to run or fail to meet functional constraints. Benchmarks such as SWE-bench (Jimenez et al., 2024) focus on real-world software engineering tasks like bug fixing in Python repositories. Due to the difference in test domains, such benchmarks cannot directly reflect a model’s ability to execute transactions in Web2/Web3 environments. Blockchain-specific benchmarks such as Solana Bench (Solana Foundation, 2025) explore the boundaries of LLM capabilities in Web3 but struggle to provide feedback on the accuracy of natural language understanding and the safety of transaction execution.
@@
We evaluate 20 models under a unified protocol. Results show substantial variance. Split scores reveal a persistent capability asymmetry between single-action precision and workflow completion. Several models achieve strong Composite performance with weaker Atomic scores. Several models fail on Composite workflows despite non-trivial Atomic performance. These patterns motivate split reporting for diagnosis.
@@
  • We report results on 20 models with split-level analysis that separates single-action precision from multi-step workflow completion.
@@ 2 Related Work
A major limitation of reference-based code evaluation is that surface overlap can reward outputs that resemble references but fail functionally. Execution-based benchmarks address this by running generated programs against tests, including HumanEval and MBPP for function synthesis (Chen et al., 2021; Austin et al., 2021), and extensions to library usage and multilingual settings (Wang et al., 2023; Khan et al., 2024; Yan and others, 2024). However, most prior benchmarks focus on stateless, sandboxed functions and do not model shared external state or irreversible actions.
@@
SWE-bench evaluates issue resolution by applying patches to real repositories and validating with test suites (Jimenez et al., 2024), emphasizing repository-scale context and cross-file edits. Separately, agent-style benchmarks study multi-step tool use and action sequencing in interactive environments (Qin et al., 2023; Liu et al., 2023). These lines are complementary, but they typically do not target blockchain-specific constraints such as transaction construction, revert conditions, strict unit handling, or protocol-imposed prerequisites.
@@
Blockchain interaction introduces shared mutable state, protocol prerequisites, and revert risk that are absent in standard code tasks. Prior blockchain benchmarks (e.g., Solana-focused transaction evaluations) (Solana Foundation, 2025) do not explicitly disentangle single-transaction precision from multi-transaction workflow completion under a unified execution and validation interface. EVM-QuestBench fills this gap by evaluating natural-language-to-transaction-script generation on EVM-compatible chains with atomic/composite splits. Tasks are specified declaratively as JSON with reusable validator components, enabling new tasks to be added with minimal engineering effort while preserving comparable execution-grounded evaluation.
@@ 3.3 Benchmark Composition
EVM-QuestBench contains 107 tasks (62 atomic, 45 composite) with a maximum total score of 10,700. Figure 2 shows split sizes and composite workflow complexity.
Figure 2: Task split and composite workflow complexity in EVM-QuestBench. (a) Atomic and composite task counts. (b) Distribution of optimal steps for composite tasks.
@@ 5.1 Leaderboard
We evaluate 20 models on EVM-QuestBench. Table 1 reports Atomic, Composite, and Total scores, where Atomic sums 62 tasks, Composite sums 45 tasks, and Total sums 107 tasks (max 10,700). Avg is computed as Total/107.
The top three models exceed 7,8007{,}800 total points. DeepSeek V3.2 achieves the strongest Composite score (4,088.34{,}088.3) while ranking mid range on Atomic (2,9152{,}915), suggesting partial separability between workflow completion and single step exactness.
# Model Atom Comp Total Avg
1 Gemini-3-Pro 4460 3757 8217 76.8
2 Kimi-K2 4400 3505 7905 73.9
3 Claude-Sonnet-4.5 4088 3781 7869 73.5
4 GPT-5.1 4145 3497 7642 71.4
5 DeepSeek-V3.2 2915 4088 7003 65.5
6 GPT-OSS-120B 2531 3682 6213 58.1
7 Grok-4-Fast 2853 3358 6211 58.0
8 Claude-Haiku-4.5 3816 2215 6031 56.4
9 Gemini-2.5-Flash 2265 3716 5981 55.9
10 GPT-5 3516 2443 $59$ 55.7
11 Qwen3-Max 2775 2698 5473 51.2
12 GPT-5.1-Codex-Mini 1573 3286 5286 49.4
13 Qwen3-235B 3580 1691 5271 49.3
14 Mimo-V2-Flash 2740 2209 $49$ 46.3
15 Qwen3-30B 2845 1887 4732 44.2
16 GLM-4.6 2771 1145 3916 36.6
17 Devstral-2512 2905 0 2905 27.2
18 Qwen3-Coder-30B 3015 0 3015 28.2
19 Qwen3-Coder 2825 0 2825 26.4
20 Qwen3-Coder-Flash 290 38 328 3.1
Table 1: Leaderboard. Atom=Atomic (62 tasks), Comp=Composite (45 tasks), Total=107 tasks. Avg=Total/107.
@@ 5.2 Atomic versus Composite Capability Separation
Figure 3 visualizes Atomic versus Composite scores. Models spread widely across the plane, indicating that single step correctness and multi step completion are only partially coupled.
We observe two characteristic asymmetries. Workflow oriented models (DeepSeek V3.2, Gemini 2.5 Flash) achieve high Composite despite weaker Atomic, suggesting stronger sequencing and end state targeting. Precision oriented models (Claude Haiku) achieve high Atomic but lag on Composite, consistent with weaknesses in multi step dependency tracking. Full quadrant assignments are provided in Appendix F.
Figure 3: Atomic score versus Composite score. Each point is a model.
@@ 5.3 Composite Execution Quality
Table 2 reports fine grained composite metrics for 13 models with complete composite evaluation logs. We analyze three dimensions: code quality (1st%), step efficiency (Eff%), and environment understanding (ethers%).
Model Score Pass K Eff 1st eth
DeepSeek-V3.2 4088 42 3.0 90.6 60.0 26.7
GPT-5 3962 40 2.2 94.9 86.7 0.0
Gemini-2.5-Flash 3876 39 3.3 85.6 57.8 0.0
Gemini-3-Pro 3703 39 2.2 97.3 71.1 4.4
GPT-OSS-120B 3682 39 3.8 85.9 51.1 4.4
Kimi-K2 3605 37 3.5 87.0 46.7 42.2
Grok-4-Fast 3358 35 4.5 76.9 24.4 60.0
GPT-5.1-Mini 3183 30 4.7 75.9 26.7 51.1
Qwen3-Max 2668 24 5.2 72.2 17.8 73.3
Claude-Haiku 2238 15 6.1 64.8 0.0 100
GLM-4.6 1145 4 6.4 62.0 0.0 100
Qwen3-Coder-Fl 38 0 6.3 56.2 0.0 86.7
Qwen3-Coder-30B 0 0 6.6 58.4 0.0 100
[…]
First try pass rate separates top models.
GPT-5 achieves 86.7% first-try pass rate with only 5 total retries, indicating high code generation quality. Gemini-3-Pro follows at 71.1%. In contrast, models with 0% first-try pass rate (Claude Haiku, GLM) require retries on every task they pass.
@@
Gemini-3-Pro leads with 97.3% efficiency despite having the same mean $K_{\text{act}}$ as GPT-5 (2.2), indicating more consistent per-task optimization. Lower tier models average $K_{\text{act}}>6$, which compounds both failure risk and score decay.
ethers% reveals environment misunderstanding.
The “ethers is not defined” failure mode indicates models that do not use the pre-provisioned ethers library correctly. Models with 0% ethers errors (GPT-5, Gemini-2.5-Flash) consistently align with top tier composite performance, while models with near-ubiquitous ethers failures (Claude Haiku, GLM) show a fundamental mismatch to the execution interface.
Several code-specialized models (e.g., qwen3-coder, devstral) report Composite scores of 0 due to early termination triggered by repeated schema/interface failures. We treat these cases as coverage artifacts and exclude them from workflow failure analysis.
@@ 6 Conclusion
We presented EVM-QuestBench, an execution grounded benchmark for natural language transaction script generation on EVM-compatible chains, instantiated on BNB Smart Chain (chain ID 56). The benchmark contains 107 tasks with Atomic and Composite splits. Evaluation executes model generated TypeScript scripts in a forked environment with snapshot isolation. Validators score outcomes after execution. Composite tasks apply a step efficiency factor.
We evaluated 20 models under a unified runner. Performance varies substantially across models. Split scores expose a persistent gap between single transaction precision and workflow completion. Several models achieve strong Composite scores with weaker Atomic scores. Several models fail on Composite tasks despite non zero Atomic performance.
EVM-QuestBench offers a standardized protocol for studying execution grounded behavior in on-chain automation. Future work can expand task coverage, extend the benchmark to additional EVM chains, and strengthen analysis using per task traces. In fact, we have already leveraged the same architecture to develop and establish a benchmark on Solana, demonstrating the portability of our atomic/composite paradigm across heterogeneous blockchain ecosystems. We also plan to incorporate richer security checks for transaction intent and side effects. Additionally, we intend to evaluate LLMs’ success rates in generating atomic and composite task definitions, which would further assess their understanding of on-chain transactions and smart contract semantics.
@@ 7 Limitations
Because EVM-QuestBench is execution grounded, scores inherit instability from RPC connectivity, fork performance, and provider availability. Snapshot isolation and composite retries reduce, but do not eliminate, this effect.
@@
We do not pin a fixed fork block height. Snapshot restore ensures comparability within a run, but cross run results can drift with mainnet state, especially for price sensitive swaps and liquidity operations. Pinning the fork height and recording the fork block number in metadata would improve reproducibility.
@@
Composite scoring applies a step efficiency decay factor $\min(1,K_{\mathrm{opt}}/K_{\mathrm{act}})$, where $K_{\mathrm{act}}$ counts execution rounds that may include retries; this can conflate planning inefficiency with execution instability. Validators also use tolerances for some tasks, which may admit occasional false positives or over penalize benign deviations. Finally, providers differ in default output budgets and formatting behavior, which can affect schema validity; prior work has noted that automatic metrics and LLM as a judge can introduce additional biases (Zheng and others, 2023; Liu and others, 2023).
Single-run evaluation.
Due to time and computational cost constraints, we evaluated each model only once rather than conducting multiple runs. This single-run protocol limits our ability to report confidence intervals or assess variance in model performance. Stochastic factors such as temperature-induced generation variability and non-deterministic execution timing may affect individual scores. Future work should incorporate multiple runs per model to provide more statistically robust comparisons.

### v2

@@ Abstract
Large language models are increasingly applied to various development scenarios. However, in on-chain transaction scenarios, even a minor error can cause irreversible loss for users. Existing evaluations often overlook execution accuracy and safety. We introduce EVM-QuestBench, an execution-grounded benchmark for natural-language transaction-script generation on EVM-compatible chains. The benchmark employs dynamic evaluation: instructions are sampled from template pools, numeric parameters are drawn from predefined intervals, and validators verify outcomes against these instantiated values. EVM-QuestBench contains 107 tasks (62 atomic, 45 composite). Its modular architecture enables rapid task development. The runner executes scripts on a forked EVM chain with snapshot isolation; composite tasks apply step-efficiency decay. We evaluate 20 models with 5 independent rounds each and find large performance gaps, with split scores revealing persistent asymmetry between single-action precision and multi-step workflow completion. Code: https://anonymous.4open.science/r/bsc_quest_bench-A9CF/.
@@ 1 Introduction
Large language models (LLMs) are increasingly being used to control software and tools (openai2023gpt4; anil2023palm2). Leveraging LLMs for code generation and blockchain transactions is becoming commonplace, but this introduces significant financial risks. Even a small error, such as an incorrect address, unit, or deadline, can result in irreversible losses.
Benchmarks for code understanding and generation need to cover a broad range of programming tasks and reasoning capabilities (chen2021evaluating; lu2021codexglue). Many evaluations still rely on lexical overlap metrics such as BLEU or CodeBLEU (papineni2002bleu; ren2020codebleu). These metrics can reward outputs that appear similar to references but fail to run or fail to meet functional constraints. Benchmarks such as SWE-bench (jimenez2024swebench) focus on real-world software engineering tasks like bug fixing in Python repositories. Due to the difference in test domains, such benchmarks cannot directly reflect a model’s ability to execute transactions in Web2/Web3 environments. Blockchain-specific benchmarks such as Solana Bench (solana2025bench) explore the boundaries of LLM capabilities in Web3 but struggle to provide feedback on the accuracy of natural language understanding and the safety of transaction execution.
@@
We evaluate 20 models under a unified protocol with 5 independent rounds per model. Results show substantial variance. Split scores reveal a persistent capability asymmetry between single-action precision and workflow completion. Several models achieve strong Composite performance with weaker Atomic scores. Several models fail on Composite workflows despite non-trivial Atomic performance. These patterns motivate split reporting for diagnosis.
@@
  • We report results on 20 models with 5-round statistical analysis, including confidence intervals and rank-order consistency, that separates single-action precision from multi-step workflow completion.
@@ 2 Related Work
A major limitation of reference-based code evaluation is that surface overlap can reward outputs that resemble references but fail functionally. Execution-based benchmarks address this by running generated programs against tests, including HumanEval and MBPP for function synthesis (chen2021evaluating; austin2021program), and extensions to library usage and multilingual settings (wang2023odex; khan2024xcodeeval; yan2024codescope). However, most prior benchmarks focus on stateless, sandboxed functions and do not model shared external state or irreversible actions.
@@
SWE-bench evaluates issue resolution by applying patches to real repositories and validating with test suites (jimenez2024swebench), emphasizing repository-scale context and cross-file edits. Separately, agent-style benchmarks study multi-step tool use and action sequencing in interactive environments (qin2023toollm; liu2023agentbench). These lines are complementary, but they typically do not target blockchain-specific constraints such as transaction construction, revert conditions, strict unit handling, or protocol-imposed prerequisites.
@@
Blockchain interaction introduces shared mutable state, protocol prerequisites, and revert risk that are absent in standard code tasks. Prior blockchain benchmarks (e.g., Solana-focused transaction evaluations) (solana2025bench) do not explicitly disentangle single-transaction precision from multi-transaction workflow completion under a unified execution and validation interface. EVM-QuestBench fills this gap by evaluating natural-language-to-transaction-script generation on EVM-compatible chains with atomic/composite splits. Tasks are specified declaratively as JSON with reusable validator components, enabling new tasks to be added with minimal engineering effort while preserving comparable execution-grounded evaluation.
@@ 3.3 Benchmark Composition
EVM-QuestBench contains 107 tasks (62 atomic, 45 composite) with a maximum total score of 10,700. Figure 2 shows the split sizes and Figure 3 shows the composite workflow complexity.
Figure 2: Task split in EVM-QuestBench: 62 atomic tasks and 45 composite tasks (107 total).
Figure 3: Distribution of optimal steps ($K_{\text{opt}}$) for composite tasks. Most workflows require 3 steps (53.3%), with a mean of $3.$
@@ 5.1 Leaderboard
We evaluate 20 models on EVM-QuestBench with 5 independent evaluation rounds per model (totaling $5\times 107\times 20=10{,}700$ task executions). Each round uses freshly sampled numeric parameters, independent LLM calls (temperature =0.7=0.7), and clean snapshot-isolated environments. Table 1 reports the mean Atomic, Composite, and Total scores across 5 rounds, along with standard deviation (SD), 95% confidence interval (CI, computed as mean $\pm\,t_{0.025,4}\times\text{SE}$ with $t=2.776$ for df =4=4), and coefficient of variation (CV%).
The top three models exceed 7,7007{,}700 mean total points. Claude-Sonnet-4.5 achieves the highest mean total (8,2368{,}236) with low variance (CV $2.1\%$). 15 out of 20 models exhibit CV% below $9\%$, demonstrating that evaluation scores are robust to the combined variance from dynamic parameter sampling, temperature-based generation, and RPC execution.
# Model Atom Comp Total SD CV% 95% CI
1 Claude-Sonnet-4.5 4278 3958 8236 174 2.1 [8020, 8451]
2 Gemini-3-Pro 4302 3561 7863 267 3.4 [7531, 8195]
3 GPT-5 4128 3646 7774 407 5.2 [7269, 8279]
4 GPT-5.1 3715 3617 7332 144 2.0 [7154, 7510]
5 Kimi-K2-Thinking 3605 3633 7238 207 2.9 [6982, 7495]
6 Gemini-2.5-Flash 3265 3768 7033 127 1.8 [6876, 7191]
7 GPT-OSS-120B 3458 3468 6926 436 6.3 [6384, 7467]
8 DeepSeek-V3.2 3071 3638 6709 384 5.7 [6232, 7186]
9 Grok-4-Fast 2980 3207 6187 224 3.6 [5909, 6465]
10 Qwen3-235B 3609 2116 5724 630 11.0 [4943, 6506]
11 GPT-5.1-Mini 2510 3190 5700 510 8.9 [5067, 6333]
12 Qwen3-Max 2775 2883 5658 229 4.1 [5374, 5943]
13 Claude-Haiku-4.5 3221 2367 5588 316 5.7 [5196, 5981]
14 Mimo-V2-Flash 3189 1723 4912 218 4.4 [4641, 5183]
15 GLM-4.6 2881 1971 4852 295 6.1 [4486, 5218]
16 Qwen3-30B 1554 2011 3565 398 11.2 [3072, 4059]
17 Devstral-2512 2892 21 2914 303 10.4 [2537, 3290]
18 Qwen3-Coder 1801 160 1961 291 14.8 [1600, 2322]
19 Qwen3-Coder-Flash 223 38 260 212 81.4 [3-3, 524]
20 Qwen3-Coder-30B 140 5 145 61 42.0 [69, 221]
Table 1: Leaderboard (avg@5). All scores are means over 5 independent rounds. Atom = Atomic (62 tasks, max 6,200), Comp = Composite (45 tasks, max 4,500), Total = 107 tasks (max 10,700). Spearman’s rank correlation between any two rounds averages $\rho=0.960$ (all $p<0.001$, range: 0.940–0.983).
@@ 5.2 Atomic versus Composite Capability Separation
Figure 4 visualizes the Atomic–Composite plane (a full ranking bar chart is in Appendix E, Figure 6). Models spread widely, indicating that single-step correctness and multi-step completion are only partially coupled.
We observe two characteristic asymmetries. Workflow-oriented models (DeepSeek-V3.2, Gemini-2.5-Flash, GPT-5.1-Mini) achieve high Composite despite weaker Atomic, suggesting stronger sequencing and end-state targeting. Precision-oriented models (Claude-Haiku-4.5, Devstral-2512) achieve high Atomic but lag on Composite, consistent with weaknesses in multi-step dependency tracking. Several code-specialized models (Qwen3-Coder-30B, Devstral-2512, Qwen3-Coder) score near zero on Composite due to repeated interface failures in multi-step workflows.
Figure 4: Atomic score versus Composite score (avg@5). Each point is a model.
@@ 5.3 Composite Execution Quality
Table 2 reports fine-grained composite metrics averaged over 5 rounds. We analyze step efficiency ($\bar{K}_{\text{act}}$, Eff%) as primary dimensions.
Model Score Pass $\bar{K}$ Eff%
Claude-Sonnet-4.5 3958 41.0 3.8 88.0
Gemini-2.5-Flash 3768 38.6 4.1 83.7
GPT-5 3646 38.0 3.6 83.8
DeepSeek-V3.2 3638 38.2 3.5 84.5
Kimi-K2-Thinking 3633 33.6 3.6 84.2
GPT-5.1 3617 38.0 4.1 80.4
Gemini-3-Pro 3561 37.6 4.3 79.7
GPT-OSS-120B 3468 35.2 3.9 82.6
Grok-4-Fast 3207 32.0 4.4 76.4
GPT-5.1-Mini 3190 31.0 4.7 73.6
Qwen3-Max 2883 25.8 4.9 71.5
Claude-Haiku-4.5 2367 13.0 5.9 60.7
Qwen3-235B 2116 11.0 5.8 62.6
Qwen3-30B 2011 8.4 6.1 58.8
GLM-4.6 1971 10.2 6.0 58.3
Mimo-V2-Flash 1723 9.0 6.2 56.9
Qwen3-Coder 160 0.4 6.7 53.3
Qwen3-Coder-Flash 38 0.2 6.8 51.1
Devstral-2512 21 0.0 6.8 50.8
Qwen3-Coder-30B 5 0.0 6.8 50.6
Table 2: Composite execution quality (avg@5, 45 tasks). Score and Pass are means over 5 rounds. $\bar{K}$ = mean $K_{\text{act}}$, Eff = step efficiency.
@@
Claude-Sonnet-4.5 leads with 88.0% efficiency and the highest composite score (3,958 avg), completing an average of 41 out of 45 tasks. DeepSeek-V3.2 achieves the lowest $\bar{K}_{\text{act}}$ (3.5) with 84.5% efficiency, indicating consistent per-task optimization. Lower-tier models average $\bar{K}_{\text{act}}>5.9$, which compounds both failure risk and score decay.
Pass count separates tiers.
Top models pass 33–41 of 45 composite tasks on average across 5 rounds. Claude-Haiku-4.5 passes only 13 tasks despite reasonable atomic performance, confirming that multi-step workflow completion requires capabilities beyond single-action precision. Four models (Devstral-2512, Qwen3-Coder-30B, Qwen3-Coder-Flash, Qwen3-Coder) pass fewer than 1 task on average, reflecting systematic interface or planning failures in multi-step workflows.
Interface compliance.
All 20 models produce syntactically valid TypeScript modules that the runner can execute on atomic tasks. However, three code-specialized models (Qwen3-Coder-30B, Devstral-2512, Qwen3-Coder) score near zero on composite tasks due to repeated schema errors such as missing ethers imports or incorrect module structure in multi-step contexts. We retain these models in the leaderboard to document the full spectrum of interface-level and capability-level failures.
@@ 6 Conclusion
We presented EVM-QuestBench, an execution-grounded benchmark for natural language transaction script generation on EVM-compatible chains, instantiated on BNB Smart Chain (chain ID 56). The benchmark contains 107 tasks with Atomic and Composite splits. Evaluation executes model generated TypeScript scripts in a forked environment with snapshot isolation. Validators score outcomes after execution. Composite tasks apply a step efficiency factor.
We evaluated 20 models under a unified runner with 5 independent evaluation rounds per model. Performance varies substantially across models. Split scores expose a persistent gap between single transaction precision and workflow completion. Several models achieve strong Composite scores with weaker Atomic scores. Several models fail on Composite tasks despite non zero Atomic performance.
EVM-QuestBench offers a standardized protocol for studying execution-grounded behavior in on-chain automation. Future work can expand task coverage, extend the benchmark to additional EVM chains, and strengthen analysis using per task traces. In fact, we have already leveraged the same architecture to develop and establish a benchmark on Solana, demonstrating the portability of our atomic/composite paradigm across heterogeneous blockchain ecosystems. We also plan to incorporate richer security checks for transaction intent and side effects. Additionally, we intend to evaluate LLMs’ success rates in generating atomic and composite task definitions, which would further assess their understanding of on-chain transactions and smart contract semantics.
@@ 7 Limitations
Because EVM-QuestBench is execution-grounded, scores inherit instability from RPC connectivity, fork performance, and provider availability. Snapshot isolation and composite retries reduce, but do not eliminate, this effect.
@@
By default, the fork block height is not pinned, and cross-run results can drift with mainnet state, especially for price-sensitive swaps and liquidity operations. Our framework supports pinning the fork at a specific block height via a configuration file; we recommend this setting for standardized comparative benchmarking. Recording the fork block number in evaluation metadata further aids reproducibility.
@@
Composite scoring applies a step efficiency decay factor $\min(1,K_{\mathrm{opt}}/K_{\mathrm{act}})$, where $K_{\mathrm{act}}$ counts execution rounds that may include retries; this can conflate planning inefficiency with execution instability. Validators also use tolerances for some tasks, which may admit occasional false positives or over penalize benign deviations. Finally, providers differ in default output budgets and formatting behavior, which can affect schema validity; prior work has noted that automatic metrics and LLM as a judge can introduce additional biases (zheng2023judging; liu2023geval).
Number of evaluation rounds.
We conduct 5 independent evaluation rounds per model and report mean scores with confidence intervals (Table 1). While 5 rounds provide reasonable statistical power for tier-level separation, some adjacent-rank differences remain statistically indistinguishable (overlapping CIs). Additional rounds would narrow confidence intervals and enable finer-grained ranking distinctions.